OpenStudy (anonymous):

subtract 8-v/v-6 - 4v-9/6-v

OpenStudy (jamesj):

get a common denominator of (v-6)

OpenStudy (anonymous):

multiply the 2nd equation by (-1) to get denominator in form of (v-6) and then work on numerators

OpenStudy (anonymous):

3v+17/v-6 ??

OpenStudy (anonymous):

nope, how about -5v-1/v-6?

OpenStudy (anonymous):

I am sorry, I mean multiply the 2nd fraction by (-1)/(-1)

OpenStudy (anonymous):

3v-1/v-6

OpenStudy (anonymous):

(8-v)/(v-6)-(4v-9)/(6-v)= (8-v)/(v-6)-(4v-9)(-1)/(6-v)(-1)= (8-v)/(v-6)+(4v-9)/(v-6)=(8-v+4v-9)/(v-6)=(3v-1)/(v-6)

OpenStudy (anonymous):

whew! finally got it lol

OpenStudy (anonymous):

good for you. It's very often will be used in your class.